Discount rate 12% Year Cash flow 1 $             675 2                 880 3    

Discount rate 12%
Year Cash flow
1 $             675
2                 880
3                 985
4              1,530
What is the present value of the cash flows?
Answers:

2468.54

2746.38

2977.66

3127.49
Answer #1
Year Cash flows Discount factor Discounted cash flows
1 675 0.892857143 602.6785714
2 880 0.797193878 701.5306122
3 985 0.711780248 701.1035441
4 1530 0.635518078 972.34266
Present value 2977.655388
Hence, the answer is $2,977.66

А 1 Year 2 1 3 2 4 3 5 4 в с Cash flows Discount factor Discounted cash flows 675 =1/1.12 =B2*C2 880 =C2/1.12 =B3*C3 985 =C3/